OpenStudy (anonymous):

an urn contains 5 red balls &5 black balls .a ball is drawn AT RANDOM ,ITS COLOUR IS NOTED AND IS RETURNED TO THE URN .MOREOVER ,2 ADDITIONAL BALLS OF THE COLOUR DRAWN ARE PuT IN THE URN AND THEN a ball is drawn at random. what is the chance that 2nd ball is red

OpenStudy (anonymous):

first figure out in what ways can you get 2nd ball as a red one. 1. BR 2. RR for first case, if you draw a black ball first and then red, so the probability will be = 5/10 x 5/12 for 2nd case prob will be = 5/10 x 7/12 add these two events and you will get the answer.